OpenStudy (anonymous):

Does the series converge or diverge?: sum of (-1)^n * ln(n)/n^(1/2)

OpenStudy (anonymous):

What's the limit as \(n\to\infty\)?

OpenStudy (paxpolaris):

\[\sum \left( -1 \right)^n {\ln \left( n \right) \over \sqrt n}\]

OpenStudy (anonymous):

the lim as n goes to infinity is \[\lim_{n \rightarrow \infty} \left| (-1)^n*\frac{ \ln(n) }{ n^(\frac{ 1 }{ 2 }) } \right|\] I started it out trying to do it with the root test, but I got stuck...

OpenStudy (anonymous):

Have you learned about the alternating series test yet?

OpenStudy (anonymous):

I have... Ill try that really quick

OpenStudy (anonymous):

HI... I have a different way. Will you like to know?. @kreimer20 ?

OpenStudy (anonymous):

yes please? I don't mind learning another way

OpenStudy (anonymous):

oh, i thought it is n^2 in the denominator.. Hmm ok you try Leibnitz's method, let me see if i can get any easy approach or not.

OpenStudy (anonymous):

Ok i got an easy method...

OpenStudy (anonymous):

lim (ln n)/n^(1/2)=0 and lim ln(n+1)/(n+1)^(1/2)=0 we can say that the series of even terms and odd terms converges to 0, And hence the series converges to 0

OpenStudy (anonymous):

@kreimer20

OpenStudy (anonymous):

I'm not too worried about finding out what it converges to.. I think my professor wants us using certain methods in this... so I did it with the alternating series and got it converging... so thanks anyway!

OpenStudy (anonymous):

@kreimer20 keep in mind that you might need to say something about conditional vs absolute convergence.